OpenStudy (anonymous):

Calculate the kinetic energy for a rolling ball. Translational kinetic energy + Rotational kinetic energy: \[K _{trans}=\dfrac{1}{2}mv ^{2}\] and \[K _{rot}=\dfrac{1}{2}I\omega^{2}\] Given infomation: \[I=\dfrac{2}{3}mR ^{2}\] and \[m=0,5kg\qquad v=1,2m/s\] My problem is, that i have no idea how to calculate the rotational kinetic energy since \[\omega\] and R aren't given.

OpenStudy (anonymous):

\[K _{rot}=\dfrac{1}{2}(\dfrac{2}{3}mR^{2})\omega ^{2}\] Which gives me \[K _{rot}=\dfrac{R^{2}w^{2}}{6}\]

OpenStudy (anonymous):

The ball is rolling so i know: \[v=R\omega\]=> \[v ^{2}=(R\omega) ^{2}\] I could then insert this into \[K _{rot}=\dfrac{R^{2}\omega ^{2}}{6}\] So i'd have: \[K _{rot}=\dfrac{v^{2}}{6}\] And i know \[v=1,2m/s\] So i'd get: \[K _{rot}=\dfrac{(1,2m/s )^{2}}{6}=0,24J\] Would that make sense ?

OpenStudy (kainui):

It appears to make sense. =)

OpenStudy (anonymous):

If you calculate the translational kinetic energy you'd get \[K _{trnas}=\dfrac{1}{2}\cdot\dfrac{1}{2}kg\cdot (1,2m/s)^{2}=0,3J\] And since the ball is rolling \[v=R\omega\] is true. Which means that the rotational energy should more or less equal to the the translational energy.

OpenStudy (anonymous):

Should be more or less equal to the translational energy*